OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. Show your work •y = x2 + 3x + 2 •y = x2 + 2x + 1 Part 2: Using complete sentences, answer the following questions about the two quadratic equations above. •Do the two quadratic equations have anything in common? If so, what? •What makes y = x2 + 3x + 2 different from y = x2 + 2x + 1?

OpenStudy (triciaal):

•y = x2 + 3x + 2 y=(x +2)(x+1) when y = 0 then x = -2, x = -1 (-2, 0) and (-1, 0)

OpenStudy (anonymous):

ok That makes sense

OpenStudy (triciaal):

y = x2 + 2x + 1 y = (x + 1)(x+1) when y = 0 x = -1 (-1, 0)

OpenStudy (triciaal):

graph 1 when x = 0 then y = 2 graph 2 when x = 0 then y = 1

OpenStudy (triciaal):

sketch the 2 graphs on the same axis describe the transformation